OpenStudy (anonymous):

Find the derivative of the function. g(x) = (2 + 3x)^4(4 + x - x^2)^7...this is awefully long to substitue...

OpenStudy (amistre64):

g(x) = (2 + 3x)^4(4 + x - x^2)^7 its product rule; power rule; and chain rule applications

OpenStudy (anonymous):

i got it up to (2+3x)^4*7(4+x-x^2)^6+(1-2x)+(4+x0x^2)^7*4(2+3x)^3*3.....

OpenStudy (amistre64):

l = (2+3x)^4 ; l' = 12(2+3x)^3 r = (4+x-x^2)^7 ; r' = 7(4+x-x^2)^6 (1-2x)

OpenStudy (anonymous):

yes..i went through that

OpenStudy (amistre64):

\[g'=(2+3x)^4.(7-14x)(4+x-x^2)^6 + (4+x-x^2)^7.12(2+3x)^3\]

OpenStudy (amistre64):

how compact are you trying to get it?

OpenStudy (anonymous):

til i cant solve no more

OpenStudy (amistre64):

\(g' =\) \( (7-14x).(4+x-x^2)^6.(2+3x)^4\) \( + 12(2+3x)^3.(4+x-x^2)^7 \) the only other thing to do is expand it all out and see what combines

OpenStudy (amistre64):

but that is usually NOT done since it is easier to work it in this form

OpenStudy (anonymous):

ok ill take it from here thanks again : )

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

\[5^{x+4}=7^{x}\]

OpenStudy (anonymous):

Can anyone tell me how to solve this to the nearest thousandth.

OpenStudy (amistre64):

odd place to post it, but maybe

OpenStudy (amistre64):

(x+4) ln(5) = x ln(7) (x+4)/x = ln(7)/ln(5) 1+4/x = ln(7)/ln(5) 4/x = [ln(7)/ln(5)] -1 x = 4/{ln(7)/ln(5)] -1} \[x = \frac{4}{\frac{ln(7)}{ln(5)}-1}\]

OpenStudy (amistre64):

19.133 i think